A projectile is fired downward with initial speed v0 in an experimental fluid and experiences an acceleration a = σ - ηv2, where σ and η are positive constants and v is the projectile speed. Determine the distance y traveled by the projectile when its speed has been reduced to 0.58v0. Also, determine the terminal velocity vf of the projectile. Evaluate for σ = 1.10 m/s2, η = 0.17 m-1, and v0 = 4.8 m/s.VYUO
Answer #1


6-1010 m/s² = 8017 in Vo = 4.8 m/s X=0.58% = 0.58 (4.8) = 2.784 m/s a=s- nr 2 vdy a 8-nut 1o-nya dva dy dy 2 کی اے 2.784 * (